OpenStudy (anonymous):

which double angle or half angle identity would you use to verify that sin2x/2sin2x=cot x

OpenStudy (anonymous):

erm, you wrote that out extremely poorly. I'm guessing you meant: \[\frac{ \sin^2(x) }{ 2\sin(2x) } = \cot(x)\] In that case, the only double angle in that formula is sin(2x) sin(2x) = 0.5*(cosx)(sinx)